Fermi National Laboratory
 

Operation Summary for August 12, 2011 - August 15, 2011

 

The Integrated Luminosity for the period from 8/8/11 to 8/15/11 was 4.11 pb-1 (inverse picobarns). NuMI reported receiving 5.94 E18 protons on target during this same period.

Friday August 12

The day shift began with Operations monitoring the Tevatron (TEV) store 8973 (70.52 E30) and the Recycler stash (311.8 E10), with the Antiproton Source (Pbar) stacking (7.3 E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBFT-1010 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8973 at 8:37 AM. The Tevatron was put into shot setup at 10:18 AM and store 8975 was colliding at 11:34 AM with an Average Initial luminosity (AIL) of 286.6 E30.

During the evening shift, Meson Test was turned off at 8:00 PM.

Saturday August 13

The midnight shift began with Operations monitoring store 8975 (73.03 E30) the stash (314.3 E10) and the Pbar stack (20.6 E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8975 at 4:01 AM. The Tevatron was put into shot setup at 4:16 AM and store 8976 was colliding at 5:13 AM with an Average Initial luminosity (AIL) of 350.5 E30. Beam was established to MTest at 7:16 AM.

During the day shift, a HVAC pipe ruptured in the Linac upper gallery spraying water onto KRF1 and KRF2. KRF2 tripped off at 2:07 PM. Experts were called to fix the ruptured pipe and dry out and repair the two klystron stations. No Linac beam was delivered the rest of the shift.

During the evening shift, repairs were completed on the ruptured HVAC pipe. Dry out and repair of the two klystrons stations continued.

Sunday August 14

The midnight shift began with Operations monitoring store 8976 (78.38 E30) the stash (242.2 E10) and the Pbar stack (8.1 E10), no beam as Linac klystron station repair continues.

During the midnight shift, Linac klystron station repair continues.

During the day shift, Linac klystron station repair continues. The Tev store was lost at 11:39 AM due to PAK1 pre-fire. Quench recovery is underway.

During the evening shift, repair of the two klystron stations was completed by 8:45 PM. Operations established beam to PBar and NuMI at 10:07 PM.

Monday August 15

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI taking beam.

During the midnight shift, a Tev quench occurred at 1:49 AM during the first HEP ramp. The Tev was turned off at 4:58 AM for an access into Tev enclosure A-E to address a stuck kautzky valve. Operations established beam to MTest at 6:30 AM.

The Plans for Monday

The plans for today are to recover and put a store in the Tev.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/15/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory